# Remarks on the Stanley depth and Hilbert depth of monomial ideals with linear quotients

### v4

@@ 1. Introduction
as $K$-vector spaces, where $m_{i}\in M$, $Z_{i}\subset\{x_{1},\ldots,x_{n}\}$ such that $m_{i}K[Z_{i}]$ is a free $K[Z_{i}]$-module. We define $\operatorname{sdepth}(\mathcal{D})=\min_{i=1}^{r}|Z_{i}|$ and
$\operatorname{sdepth}(M)=\max\{\operatorname{sdepth}(\mathcal{D})|\;\mathcal{D}\text{ is a Stanley decomposition of }M\}.$
@@
Given a monomial ideal $I\subset S$, Soleyman Jahan [14] noted that $I$ satisfies the Stanley inequality, i.e.
@@
In Theorem 2.4, we prove that if $I\subset S$ is a monomial ideal with linear quotients with $\operatorname{depth}(S/I)=n-2$, then $\operatorname{sdepth}(S/I)=n-2$. In Theorem 2.6, we prove that if $I\subset S$ is a monomial ideal with linear quotients which has a Stanley decomposition which satisfies certain conditions, then $\operatorname{sdepth}(S/I)\geq\operatorname{depth}(S/I)$. Also, we conjecture that for any monomial ideal $I\subset S$ with linear quotients, there is a variable $x_{i}$ such that $\operatorname{depth}(S/(I,x_{i}))=\operatorname{depth}(S/I)$ and $\operatorname{sdepth}(S/(I,x_{i}))\leq\operatorname{sdepth}(S/I)$. In Theorem 2.12 we prove that if this conjecture is true, then $\operatorname{sdepth}(S/I)\geq\operatorname{depth}(S/I)$, for any monomial ideal $I\subset S$ with linear quotients.
@@
$\operatorname{hdepth}(M)=\max\left\{r\;:\;\Large\begin{subarray}{c}\text{ There exists a f.g. graded }S\text{-module }N\\ \text{ with }H_{M}(t)=H_{N}(t)\text{ and }\operatorname{depth}(N)=r\end{subarray}\normalsize\right\}.$
@@
Let $0\subset I\subsetneq J\subset S$ be two squarefree monomial ideals. For any $0\leq j\leq n$, we let $\alpha_{j}(J/I)$ to be the number of squarefree monomials $u\in S$ of degree $j$ such that $u\in J\setminus I$. Also, for $0\leq k\leq d\leq n$, let
@@
Note that, from (1.1), one can easily see that
$\operatorname{hdepth}(J/I)\leq\max\{k\;:\;\alpha_{k}(J/I)>0\}.$ (1.2)
Also, we will make use of the fact that
@@
In Section 3 of our paper we study the Hilbert depth of $S/I$, where $I$ is a squarefree monomial ideal with linear quotients. In Proposition 3.2 we compute the numbers $\beta_{k}^{d}(I)$’s and $\beta_{k}^{d}(S/I)$’s. In Corollary 3.3, we express this numbers in terms of hypergeometric functions, thus showing the difficulty in finding explicit formulas for $\operatorname{hdepth}(I)$ and $\operatorname{hdepth}(S/I)$.
@@
This result makes use of the fact that we can assume $\deg(u_{1})\geq\cdots\geq\deg(u_{m})$; see [15].
@@ 2. Main results
According to [12, Corollary 2.7], the projective dimension of $S/I$ is
@@
Let $I\subset S$ be a monomial ideal and $u\in S$ a monomial. Then:
@@
    $\operatorname{depth}(S/(I:u))\geq\operatorname{depth}(S/I)$. ([11, Corollary 1.3])
@@
Let $0\to U\to M\to N\to 0$ be a short exact sequence of $\mathbb{Z}^{n}$-graded $S$-modules. Then:
  1. (1)
    $\operatorname{depth}(M)\geq\min\{\operatorname{depth}(U),\operatorname{depth}(N)\}$. (Depth Lemma)
  2. (2)
    $\operatorname{sdepth}(M)\geq\min\{\operatorname{sdepth}(U),\operatorname{sdepth}(N)\}$. ([11, Lemma 2.2])
Proposition 2.3.
Let $I\subset S$ be a proper monomial ideal with linear quotients with $\operatorname{depth}(S/I)=n-s-1$, where $10\leq s\leq n-$ 1. Then there exists a subset $\tau\subset[n]$ with $|\tau|=s$ and a monomial $u\in G(I)$, such that $I+(x_{i}\;:\;i\in\tau)=(u)+(x_{i}\;:\;i\in\tau)$.
@@
If $I=(u)$ is principal, that is $s=0$, then there is nothing to prove. Assume $s\geq 1$. Since $I$ has linear quotients, we can assume that $G(I)=\{u_{1},\ldots,u_{m}\}$ such that $((u_{1},\ldots,u_{j-1}):u_{j})$ is generated by variables, for every $2\leq j\leq m$. We consider the decomposition (2.1), that is
@@
where $Z_{1}=\{x_{1},\ldots,x_{n}\}$ and $Z_{j}$ is generated by the variables which do not belong to $((u_{1},\ldots,u_{j-1}):u_{j})$, for $2\leq j\leq m$. From (2.2), it follows that $|Z_{j}|\geq s$ for $2\leq j\leq m$. Note that, for any $2\leq j\leq m$, we have
$(u_{1},\ldots,u_{j-1})\cap u_{j}K[Z_{j}]=\{0\}$ (2.3)
We assume, by contradiction, that for any $\tau\subset[n]$ with $|\tau|=s$, there exists $k_{\tau}\neq\ell_{\tau}\in[m]$ such that $u_{k_{\tau}},u_{\ell_{\tau}}\notin(x_{i}\;:\;i\in\tau)$, that is $u_{k_{\tau}},u_{\ell_{\tau}}\in K[x_{i}\;:\;i\notin\tau]$. We claim that there exists $i=i(\tau)\in\tau$ such that $x_{i}\in Z_{k_{\tau}}$ or $x_{i}\in Z_{\ell_{\tau}}$. Indeed, otherwise, since $|Z_{k_{\tau}}|,|Z_{k_{\ell}}|\geq s$ we would have $Z_{k_{\tau}}=Z_{k_{\ell}}=\{x_{i}\;:\;i\notin\tau\}$ and hence
$\operatorname{gcd}(u_{k_{\tau}},u_{\ell_{\tau}})\in u_{k_{\tau}}K[Z_{k_{\tau}}]\cap u_{\ell_{\tau}}K[Z_{\ell_{\tau}}],$
a contradiction. Without any loss of generality, we may assume that $x_{i}\in Z_{k_{\tau}}$.
Let $\sigma=\{n-s+1,\ldots,n\}$. By reordering the variables, we can assume that $k_{\sigma}\geq k_{\tau}$ for any $\tau\subset[n]$ with $|\tau|=s$. Note that $u_{k_{\sigma}}=x_{1}^{a_{1}}\cdots x_{s}^{a_{n-s}}$. By the above argument, there exists $i\geq n-s+1$ such that $x_{i}\in Z_{k_{\sigma}}$. Let $A\subset[n]$ with $|A|=n-s$ such that $i\in A$ and $\{x_{\ell}\;:\;\ell\in A\}\subset Z_{k_{\sigma}}$. It follows that $\tau=[n]\setminus A\neq\sigma$. On the other hand,
$\operatorname{gcd}(u_{k_{\tau}},u_{k_{\sigma}})\in u_{k_{\tau}}S\cap u_{k_{\sigma}}K[Z_{k_{\sigma}}],$
which contradicts (2.3) for $j=k_{\sigma}$. Hence, the proof is complete. ∎
Note that, a monomial ideal $I\subset S$ is principal if and only if $\operatorname{depth}(S/I)=n-1$ if and only if $\operatorname{sdepth}(S/I)=n-1$.
@@
We assume that $G(I)=\{u_{1},\ldots,u_{m}\}$. We use induction on $m$ and $d=\sum_{j=1}^{m}\deg(u_{i})$. If $m=2$, then from [4, Proposition 1.6] it follows that $\operatorname{sdepth}(S/I)=n-2$. If $d=2$, then $I$ is generated by two variables and there is nothing to prove.
Assume $m>2$ and $d>2$. Acording to Proposition 2.3, there exists $i\in[n]$ and $u_{k}\in G(I)$ such that $(I,x_{i})=(u_{k},x_{i})$. Since $(I,x_{i})=(u_{k},x_{i})$, from [4, Proposition 1.2] it follows that $\operatorname{sdepth}(S/(I,x_{i}))\geq n-2$. If $(I:x_{i})$ is principal, then $\operatorname{sdepth}(S/(I:x_{i}))=n-1$. Note that, at least one of $u_{j}$’s must be disivible with $x_{i}$, otherwise we obtain
$G(I,x_{i})=\{u_{1},\ldots,u_{m},x_{i}\},$
a contradiction with the fact that $(I,x_{i})=(u_{k},x_{i})$ and $m\geq 2$. It follows that
$d^{\prime}:=\sum_{u\in G(I:x_{i})}\deg u<d,$
@@
From Proposition $1.2(2)$ and the short exact sequence
@@
such that there exists $i_{0}$ with $Z_{i_{0}}=\{x_{m+1},\ldots,x_{n}\}$ and $v_{i_{0}}|u$. Then:
@@
$S/(I,u)\cong\left(\bigoplus_{i\neq i_{0}}v_{i}K[Z_{i}]\right)\oplus\frac{v_{i_{0}}K[Z_{i_{0}}]}{uK[Z_{i_{0}}]}\cong\left(\bigoplus_{i\neq i_{0}}v_{i}K[Z_{i}]\right)\oplus v_{i_{0}}\frac{K[x_{m+1},\ldots,x_{n}]}{w_{0}K[x_{m+1},\ldots,x_{n}]},$ (2.6)
where $w_{0}=\frac{u}{v_{i_{0}}}$. On the other hand, $\operatorname{sdepth}\left(\frac{K[x_{m+1},\ldots,x_{n}]}{w_{0}K[x_{m+1},\ldots,x_{n}]}\right)=n-m-1$. Hence (2.6) yields the required conclusion. ∎
@@
Let $I\subset S$ be a monomial ideal with linear quotients, $G(I)=\{u_{1},\ldots,u_{m}\}$, $I_{j}=(u_{1},\ldots,u_{j})$ for $1\leq j\leq m$, such that $(I_{j-1}:u_{j})=(\{x_{1},\ldots,x_{n}\}\setminus Z_{j})$, where $Z_{j}\subset\{x_{1},\ldots,x_{n}\}$, for all $2\leq j\leq m$.
We assume that for any $2\leq j\leq m$, there exists a Stanley decomposition $\mathcal{D}_{j-1}$ of $S/I_{j-1}$ such that $\operatorname{sdepth}(\mathcal{D}_{j-1})\geq\operatorname{depth}(S/I_{j-1})$ and there exists a Stanley subspace $w_{j-1}K[W_{j-1}]$ of $\mathcal{D}_{j-1}$ with $w_{j-1}|u_{j}$ and $W_{j-1}=Z_{j}$.
@@
$\operatorname{sdepth}(S/I_{j})=\operatorname{sdepth}(S/(I_{j-1},u_{j}))\geq\min\{\operatorname{sdepth}(\mathcal{D}_{j-1}),n-n_{j}-1\}\}\geq$
$\geq\min\{\operatorname{depth}(S/I_{j-1}),n-n_{j}-1\},\text{ for all }2\leq j\leq m,$ (2.7)
@@
$\operatorname{depth}(S/I)=\min_{j=1}^{m}\{n-n_{j}-1\}.$ (2.8)
Since $\operatorname{sdepth}(S/I_{1})=\operatorname{depth}(S/I_{1})=n-1$, from (2.7) and (2.8) we get the required conclusion. ∎
@@
$I=u_{1}K[Z_{1}]\oplus u_{2}K[Z_{2}]\oplus u_{3}K[Z_{3}]=x_{1}^{2}K[x_{1},x_{2},x_{3},x_{4}]\oplus x_{1}x_{2}^{2}K[x_{2},x_{3},x_{4}]\oplus x_{1}x_{2}x_{3}K[x_{3},x_{4}].$
@@
of $S/I_{1}$ with $\operatorname{sdepth}(\mathcal{D}_{1})=\operatorname{sdepth}(S/I_{1})=2$. Let $w_{1}=x_{1}$ and $W_{1}=\{x_{2},x_{3},x_{4}\}$. Clearly, $W_{1}=Z_{2}$ and $w_{1}\mid u_{2}$. As in the proof of Lemma 2.5, we obtain the Stanley decomposition
$\mathcal{D}_{2}:\;S/I_{2}=K[x_{2},x_{3},x_{4}]\oplus\frac{x_{1}K[x_{2},x_{3},x_{4}]}{x_{1}x_{2}^{2}K[x_{2},x_{3},x_{4}]}=K[x_{2},x_{3},x_{4}]\oplus x_{1}K[x_{3},x_{4}]\oplus x_{1}x_{2}K[x_{3},x_{4}].$
Let $w_{2}=x_{1}x_{2}$ and $W_{2}=\{x_{3}\}$. Clearly, $W_{2}=Z_{3}$ and $w_{2}\mid u_{3}$. Hence, according to Theorem 2.6, $\operatorname{sdepth}(S/I)\geq\operatorname{depth}(S/I)$. In fact, we have that $\operatorname{sdepth}(S/I)\geq\operatorname{depth}(S/I)=0$ and
@@
is a Stanley decomposition of $S/I$ with $\operatorname{sdepth}(\mathcal{D})=\operatorname{sdepth}(S/I)=1$.

### v5

@@ 1. Introduction
as $K$-vector spaces, where $m_{i}\in M$ are homogeneous, $Z_{i}\subset\{x_{1},\ldots,x_{n}\}$ such that $m_{i}K[Z_{i}]$ is a free $K[Z_{i}]$-module; $m_{i}K[Z_{i}]$ is called a Stanley subspace of $M$. We define $\operatorname{sdepth}(\mathcal{D})=\min_{i=1}^{r}|Z_{i}|$ and
$\operatorname{sdepth}(M)=\max\{\operatorname{sdepth}(\mathcal{D})\;:\;\mathcal{D}\text{ is a Stanley decomposition of }M\}.$
@@
Given a monomial ideal with linear quotients $I\subset S$, Soleyman Jahan [11] noted that $I$ satisfies the Stanley inequality, i.e.
@@
In Theorem 2.4, we prove that if $I\subset S$ is a monomial ideal with linear quotients with $\operatorname{depth}(S/I)=n-2$, then $\operatorname{sdepth}(S/I)=n-2$. In Theorem 2.6, we prove that if $I\subset S$ is a monomial ideal with linear quotients which has a Stanley decomposition which satisfies certain conditions, then $\operatorname{sdepth}(S/I)\geq\operatorname{depth}(S/I)$. Also, we conjecture that for any monomial ideal $I\subset S$ with linear quotients, there is a variable $x_{i}$ such that $\operatorname{depth}(S/(I,x_{i}))\geq\operatorname{depth}(S/I)$ and $\operatorname{sdepth}(S/(I,x_{i}))\leq\operatorname{sdepth}(S/I)$. In Theorem 2.12 we prove that if this conjecture is true, then $\operatorname{sdepth}(S/I)\geq\operatorname{depth}(S/I)$, for any monomial ideal $I\subset S$ with linear quotients.
@@
$\operatorname{hdepth}(M)=\max\left\{r\;:\;\begin{subarray}{c}\text{ There exists a f.g. graded }S\text{-module }N\\ \text{ with }H_{M}(t)=H_{N}(t)\text{ and }\operatorname{depth}(N)=r\end{subarray}\right\}.$
@@
Let $0\subset I\subsetneq J\subset S$ be two squarefree monomial ideals. For any $0\leq j\leq n$, we let $\alpha_{j}(J/I)$ to be the number of squarefree monomials $u\in S$ of degree $j$ such that $u\in J\setminus I$. (In particular, $\alpha_{j}(I)$ is the number of squarefree monomials of degree $j$ which belong to $I$ and $\alpha_{j}(S/I)=\binom{n}{j}-\alpha_{j}(I)$ is the number of squarefree monomials of degree $j$ which do not belong to $I$.)
Also, for $0\leq k\leq d\leq n$, we let
@@
(In particular, $\beta_{k}^{d}(S/I)=\sum_{j=0}^{k}(-1)^{k-j}\binom{d-j}{k-j}\alpha_{j}(S/I)$ and $\beta_{k}^{d}(I)=\sum_{j=0}^{k}(-1)^{k-j}\binom{d-j}{k-j}\alpha_{j}(I)$.)
From (1.1), using an inversion formula, it follows that
$\alpha_{k}(J/I)=\sum_{j=0}^{k}\binom{d-j}{k-j}\beta_{j}^{d}(J/I)\text{ for all}0\leq k\leq d\leq n.$ (1.2)
@@
If $I\subset S$ is a proper squarefree monomial ideal, we claim that
$\operatorname{hdepth}(S/I)\leq\max\{k\;:\;\alpha_{k}(S/I)>0\}.$ (1.3)
Note that $\alpha_{n}(S/I)=0$, since $x_{1}\cdots x_{n}\in I$, and thus $m:=\max\{k\;:\;\alpha_{k}(S/I)>0\}<n$. From (1.3) it follows that
$\alpha_{m+1}(S/I)=\sum_{j=0}^{m+1}\beta^{m+1}_{j}(S/I).$
Since $I\neq S$ it follows that $1\notin I$ and thus $\beta^{m+1}_{0}(S/I)=\alpha_{0}(S/I)=1$. The above identity implies that there exists some $11\leq k\leq m+$ 1 with $\beta_{k}^{m+1}(S/I)<0$ and therefore $\operatorname{hdepth}(S/I)\leq m$, as required.
Also, we will make use of the well known fact that
@@
In Section 3 of our paper we study the Hilbert depth of $S/I$, where $I$ is a squarefree monomial ideal with linear quotients. In Proposition 3.2 we compute the numbers $\beta_{k}^{d}(I)$’s and $\beta_{k}^{d}(S/I)$’s. In Corollary 3.3, we express these numbers in combinatorial terms, thus showing the difficulty in finding explicit formulas for $\operatorname{hdepth}(I)$ and $\operatorname{hdepth}(S/I)$.
@@ 2. Main results
According to [10, Corollary 2.7], the projective dimension of $S/I$ is
@@
Let $I\subset S$ be a monomial ideal and $u\in S\setminus I$ a monomial. Then:
@@
    $\operatorname{depth}(S/(I:u))\geq\operatorname{depth}(S/I)$. ([9, Corollary 1.3])
@@
Let $0\to U\to M\to N\to 0$ be a short exact sequence of finitely generated $\mathbb{Z}^{n}$-graded $S$-modules. Then $\operatorname{sdepth}(M)\geq\min\{\operatorname{sdepth}(U),\operatorname{sdepth}(N)\}$. ([9, Lemma 2.2])
Note that, a proper monomial ideal $I\subset S$ is principal if and only if $\operatorname{depth}(S/I)=n-1$ if and only if $\operatorname{sdepth}(S/I)=n-1$.
Lemma 2.3.
Let $I\subset S$ be a monomial ideal with linear quotients with $\operatorname{depth}(S/I)=n-2$. Then there exists some $i\in[n]$ and a monomial $u\in G(I)$ such that $(I,x_{i})=(u,x_{i})$ and $(I:x_{i})$ has linear quotients.
@@
If $n=2$ and $I=S$ then $u=1\in G(I)$ and the assertion is obvious. Hence, we may assume that $I$ is proper.
First, note that $I$ is not principal. Since $I$ has linear quotients, we can assume that $G(I)=\{u_{1},\ldots,u_{m}\}$ such that $((u_{1},\ldots,u_{j-1}):u_{j})$ is generated by variables, for every $2\leq j\leq m$. We consider the decomposition (2.1), that is
@@
where $Z_{1}=\{x_{1},\ldots,x_{n}\}$ and $Z_{j}$ is the set of variables which do not belong to $((u_{1},\ldots,u_{j-1}):u_{j})$, for $2\leq j\leq m$. From (2.2), it follows that $|Z_{j}|=n-1$ for $2\leq j\leq m$. Since $((u_{1},\ldots,u_{j-1}):u_{j})=(x_{i}\;:\;x_{i}\notin Z_{j})$, it follows that for any $2\leq j\leq m$ we have
$(u_{1},\ldots,u_{j-1})\cap u_{j}K[Z_{j}]=\{0\}.$ (2.3)
We assume, by contradiction, that for any $i\in[n]$ there exists $k_{i}>\ell_{i}\in[m]$ such that $x_{i}\nmid u_{\ell_{i}}$. We claim that $x_{i}\in Z_{k_{i}}$. Indeed, otherwise $Z_{k_{i}}=\{x_{1},\ldots,x_{n}\}\setminus\{x_{i}\}$ and therefore $u_{k_{i}}u_{\ell_{i}}\in u_{\ell_{i}}S\cap u_{k_{i}}K[Z_{k_{i}}]$, a contradiction to (2.3) for $j=k_{i}$.
Without any loss of generality, we can assume $k_{n}=\max\{k_{i}\;:\;i\in[n]\}$. Since $x_{n}\in Z_{k_{n}}$, it follows that $Z_{k_{n}}=\{x_{1},\ldots,x_{n}\}\setminus\{x_{t}\}$ for some $t\leq n-1$. Since $x_{t}\in Z_{k_{t}}$ it follows that $k_{t}<k_{n}$ and, moreover, $x_{t}\nmid u_{k_{t}}$, that is $u_{k_{t}}\in K[Z_{k_{n}}]$. Therefore $u_{k_{t}}u_{k_{n}}\in u_{k_{t}}S\cap u_{k_{n}}K[Z_{k_{n}}]$, a contradiction to (2.3) for $j=k_{n}$.
Thus, that there exists $i\in[n]$ such that for any $k_{i}>\ell_{i}\in[m]$, $x_{i}\mid u_{\ell_{i}}$. It implies that $x_{i}\mid u_{j}$ for $j=1,\ldots,m-1$. It follows that $(I:x_{i})=(u^{\prime}_{1},\ldots,u^{\prime}_{m})$ where $u^{\prime}_{j}=u_{j}/x_{i}$ for $j=1,\ldots,m-1$ and $u^{\prime}_{m}=u_{m}$ if $x_{i}\nmid u_{m}$ and $u^{\prime}_{m}=u_{m}/x_{i}$ if $x_{i}\mid u_{m}$. It is clear that $\{u^{\prime}_{1},\ldots,u^{\prime}_{m-1}\}\subset G(I:x_{i})$ and
$((u^{\prime}_{1},\ldots,u^{\prime}_{j-1}):u^{\prime}_{j})=((u_{1},\ldots,u_{j-1}):u_{j})\text{ for all }2\leq j\leq m-1.$ (2.4)
We have $u^{\prime}_{m}\in(u_{m}:x_{i})\subseteq(I:x_{i})$. If $u^{\prime}_{m}\notin G(I:x_{i})$ then $G(I:x_{i})=\{u^{\prime}_{1},\ldots,u^{\prime}_{m-1}\}$ and, from (2.4), it follows that $(I:x_{i})$ has linear quotients. On the other hand, assume $u^{\prime}_{m}\in G(I:x_{i})$ then we claim that $x_{i}\mid u_{m}$, hence $((u^{\prime}_{1},\ldots,u^{\prime}_{m-1}):u^{\prime}_{m})=((u_{1},\ldots,u_{m-1}):u_{m})$ and, again, from (2.4), it follows that $(I:x_{i})$ has linear quotients. Indeed, otherwise, $u^{\prime}_{m}=u_{m}\in G(I:x_{i})$ and $G(I:x_{i})=\{u^{\prime}_{1},\ldots,u^{\prime}_{m-1},u_{m}\}$, then for $11\leq j\leq m-$ 1 there exists $\ell_{j}\in[m]\setminus[i]$ such that $x_{\ell_{j}}\mid u^{\prime}_{j}$ and $x_{\ell_{j}}\nmid u_{m}$, and thus $x_{\ell_{j}},x_{i}\in\operatorname{supp}(u_{i})\setminus\operatorname{supp}(u_{m})$, which contradicts that $((u_{1},\ldots,u_{m-1}):u_{m})$ is generated by variables. ∎
@@
If $n=2$ then $S/I=0$ and there is nothing to prove, so we may assume $n\geq 3$ and $I$ is proper with $G(I)=\{u_{1},\ldots,u_{m}\}$ for some $m\geq 2$. We use induction on $m$ and $d=\sum_{j=1}^{m}\deg(u_{i})$. If $m=2$, then from [4, Proposition 1.6] it follows that $\operatorname{sdepth}(S/I)=n-2$. If $d=2$, then $I$ is generated by two variables and there is nothing to prove.
Assume $m>2$ and $d>2$. According to Lemma 2.3, there exist $i\in[n]$ such that $(I,x_{i})=(u_{m},x_{i})$. Since $(I,x_{i})=(u_{m},x_{i})$, from [4, Proposition 1.2] it follows that $\operatorname{sdepth}(S/(I,x_{i}))\geq n-2$. If $(I:x_{i})$ is principal, then $\operatorname{sdepth}(S/(I:x_{i}))=\operatorname{depth}(S/(I:x_{i}))=n-1$.
Assume that $(I:x_{i})$ is not principal. We have that $\operatorname{depth}(S/(I:x_{i}))\leq n-2$. On the other hand, by Proposition 2.1(1) we have $\operatorname{depth}(S/(I:x_{i}))\geq\operatorname{depth}(S/I)=n-2$ and thus $\operatorname{depth}(S/(I:x_{i}))=n-2$. From the proof of Lemma 2.3, we have $G(I:x_{i})\subset\{u_{1}/x_{i},\ldots,u_{m-1}/x_{i},u_{m}\}$. It follows that
$d^{\prime}:=\sum_{u\in G(I:x_{i})}\deg(u)<d,$
@@
From Proposition 2.2 and the short exact sequence
@@
such that there exists $i_{0}$ with $Z_{i_{0}}=\{x_{m+1},\ldots,x_{n}\}$ and $v_{i_{0}}\mid u$. Then:
@@
$S/(I,u)\cong\left(\bigoplus_{i\neq i_{0}}v_{i}K[Z_{i}]\right)\oplus\frac{v_{i_{0}}K[Z_{i_{0}}]}{uK[Z_{i_{0}}]}\cong\left(\bigoplus_{i\neq i_{0}}v_{i}K[Z_{i}]\right)\oplus\frac{K[x_{m+1},\ldots,x_{n}]}{w_{0}K[x_{m+1},\ldots,x_{n}]},$ (2.7)
where $w_{0}=\frac{u}{v_{i_{0}}}$. On the other hand, $\operatorname{sdepth}\left(\frac{K[x_{m+1},\ldots,x_{n}]}{w_{0}K[x_{m+1},\ldots,x_{n}]}\right)=n-m-1$. Hence (2.7) and Proposition 2.2 yields the required conclusion. ∎
@@
Let $I\subset S$ be a monomial ideal with linear quotients, $G(I)=\{u_{1},\ldots,u_{m}\}$. Let $I_{j}=(u_{1},\ldots,u_{j})$ for $1\leq j\leq m$, such that $(I_{j-1}:u_{j})=(\{x_{1},\ldots,x_{n}\}\setminus Z_{j})$, where $Z_{j}\subset\{x_{1},\ldots,x_{n}\}$, for all $2\leq j\leq m$.
We assume that for any $2\leq j\leq m$, there exists a Stanley decomposition $\mathcal{D}_{j-1}$ of $S/I_{j-1}$ such that $\operatorname{sdepth}(\mathcal{D}_{j-1})=\operatorname{sdepth}(S/I_{j-1})$ and there exists a Stanley subspace $w_{j-1}K[W_{j-1}]$ of $\mathcal{D}_{j-1}$ with $w_{j-1}\mid u_{j}$ and $W_{j-1}=Z_{j}$.
@@
$\operatorname{sdepth}(S/I_{j})=\operatorname{sdepth}(S/(I_{j-1},u_{j}))\geq\min\{\operatorname{sdepth}(\mathcal{D}_{j-1}),n-n_{i}-1\}$
$=\min\{\operatorname{sdepth}(S/I_{j-1}),n-n_{j}-1\}\},\text{ for all }2\leq j\leq m,$ (2.8)
@@
$\operatorname{depth}(S/I)=\min_{j=2}^{m}\{n-n_{j}-1\}.$ (2.9)
Since $\operatorname{sdepth}(S/I_{1})=\operatorname{depth}(S/I_{1})=n-1$, by applying repeatedly (2.8) we deduce that
$\operatorname{sdepth}(S/I)=\operatorname{sdepth}(S/I_{m})\geq\min_{j=2}^{m}\{n-n_{j}-1\}.$
Hence, from (2.9) we get the required conclusion. ∎
@@
$I=u_{1}K[Z_{1}]\oplus u_{2}K[Z_{2}]\oplus u_{3}K[Z_{3}]=x_{1}^{2}K[x_{1},x_{2},x_{3},x_{4}]\oplus x_{1}x_{2}^{2}K[x_{2},x_{3},x_{4}]\oplus x_{1}x_{2}x_{3}^{2}K[x_{3},x_{4}].$
@@
of $S/I_{1}$ with $\operatorname{sdepth}(\mathcal{D}_{1})=\operatorname{sdepth}(S/I_{1})=3$. Let $w_{1}=x_{1}$ and $W_{1}=\{x_{2},x_{3},x_{4}\}$. Clearly, $W_{1}=Z_{2}$ and $w_{1}\mid u_{2}$. As in the proof of Lemma 2.5, we obtain the Stanley decomposition
$\mathcal{D}_{2}:\;S/I_{2}=K[x_{2},x_{3},x_{4}]\oplus\frac{x_{1}K[x_{2},x_{3},x_{4}]}{x_{1}x_{2}^{2}K[x_{2},x_{3},x_{4}]}=K[x_{2},x_{3},x_{4}]\oplus x_{1}K[x_{3},x_{4}]\oplus x_{1}x_{2}K[x_{3},x_{4}]$
of $S/I_{2}$ with $\operatorname{sdepth}(\mathcal{D}_{2})=\operatorname{sdepth}(S/I_{2})=2$.
Let $w_{2}=x_{1}x_{2}$ and $W_{2}=\{x_{3},x_{4}\}$. Clearly, $W_{2}=Z_{3}$ and $w_{2}\mid u_{3}$. Hence, according to Theorem 2.6, $\operatorname{sdepth}(S/I)\geq\operatorname{depth}(S/I)=1$. Note that
@@
is a Stanley decomposition of $S/I$ with $\operatorname{sdepth}(\mathcal{D})=1$ and thus $\operatorname{sdepth}(S/I)\geq 1$.
On the other hand, since $(x_{1},x_{2},x_{3})$ is an associated prime to $S/I$, it follows that $\operatorname{sdepth}(S/I)\leq 1$ and thus $\operatorname{sdepth}(S/I)=1$. Finally, note that
$\operatorname{depth}(S/I_{2})=2,\;(I_{2},x_{1})=(x_{1})\text{ and }(I_{2}:x_{1})=(x_{1},x_{2}^{2}).$
In particular, we have $\operatorname{sdepth}(S/(I_{2}:x_{1}))=\operatorname{depth}(S/(I_{2}:x_{1}))=2$, while $\operatorname{sdepth}(S/(I_{2},x_{1}))=\operatorname{depth}(S/(I_{2},x_{1}))=3$.
@@
Lemma 2.9.
Let $I\subset S$ be a monomial ideal with linear quotients and $x_{i}$ a variable. Then $(x_{i},I)$ has linear quotients. Moreover, if $S^{\prime}=K[x_{1},\ldots,x_{i-1},x_{i+1},\ldots,x_{n}]$, then $(x_{i},I)=(x_{i},J)$, where $J\subset S^{\prime}$ is a monomial ideal with linear quotients.
Proof.
$((x_{i},u_{j_{1}},\ldots,u_{j_{k-1}}):u_{j_{k}})=(x_{i},\bar{Z}_{j_{k}}).$ (2.10)
Indeed, since $((u_{1},\ldots,u_{j_{k}-1}):u_{j_{k}})=(\bar{Z}_{j_{k}})$ and $x_{i}u_{j_{k}}\in(x_{i},u_{j_{1}},\ldots,u_{j_{k-1}})$ it follows that $(x_{i},\bar{Z}_{j_{k}})\subset((x_{i},u_{j_{1}},\ldots,u_{j_{k-1}}):u_{j_{k}})$. Conversely, assume that $v\in S$ is a monomial with $vu_{j_{k}}\in(x_{i},u_{j_{1}},\ldots,u_{j_{k-1}})=(x_{i},u_{1},\ldots,u_{j_{k}-1})$. If $x_{i}\nmid v$, then $vu_{j_{k}}\in(u_{1},\ldots,u_{j_{k}-1})$, hence $v\in(\bar{Z}_{j_{k}})$. If $x_{i}\mid v$, then $v\in(x_{i},\bar{Z}_{j_{k}})$. Hence the claim (2.10) is true and therefore $(x_{i},I)$ has linear quotients. Now, let $J=(u_{j_{1}},\ldots,u_{j_{p}})$. For any $2\leq k\leq p$, we have that